GMOS-N data taken between the nights of 31 August 2020 and 8 September 2020 show a new flat field feature from dust in the optics, as marked in the image below. A preliminary cleaning of the accessible GMOS optics performed during daytime on 3 September 2020 was not successful in removing the new dust feature and the overall effects were found to be negligible.
GPI is now warming up since Monday the 16th due to the unplanned shutdown due to the Covid-19 outbreak. The instrument is now safe in a shutdown state on the telescope.
As Gemini South is closed for operations from last Monday due to the Covid-19 outbreak (see https://www.gemini.edu/news/gemini-south-shutdown-until-further-notice), GMOS-S has been warmed up and switched off. This shutdown will last until further notice
Zorro has now an H-alpha filter! Addittionally, the same new control software has also been implemented. The first Zorro run of the semester is ongoing.
F2 will be off the telescope until March 27th, due to work on the instrument cold heads.
GPI is not offered in Semester 2020B as the instrument will leave GS in early 20B for upgrades and later commissioning at GN.
In January, comprehensive testing of the spectrograph subsystems was completed yielding excellent results, e.g.
The Gemini North Adaptive Optics system, Altair, is currently offline due to problems with its deformable mirror (DM) electronics. Issues with other mechanisms on the optical bench have been resolved, however, we are still working on finding the root cause of the DM electronics failure. We are making progress but still need to test a number of components. We will have more updates by the end of February. Please email your contact scientist if you have any questions or concerns about the potential impact on your program.
We are using a new version of the 'Alopeke control software. Be aware that the FITS headers have changed slightly.
After the short thermal cycle on Dec 18, 2019 to replace the coldhead's displacer, the charge smearing in GMOS-S CCD1 has re-appeared as shown in the screenshot of the N&S dark from April 2019. Any programs using the N&S mode are affected and the programs set up with IFU-2 or IFU-R observations are possibly affected by this as well. PIs of those programs are contacted to assess the impact on their data quality and the programs are set on hold. A controlled thermal cycle is scheduled to resolve the issue at the end of January.